OpenStudy (anonymous):

Needing help please - why can't I use L'Hopitals rule for lim x-> infinity x^1/x

OpenStudy (zarkon):

\[x^{1/x}=e^\frac{\ln(x)}{x}\]

OpenStudy (anonymous):

What does your equation look like? x^1/x?

OpenStudy (anonymous):

@zordoloom yes

OpenStudy (zzr0ck3r):

use what zarkon showed you. x^(1/x) = e^(ln(x)/x) take the derivative of the ratio in the exponent(top and bottom) e^((1/x)/1) = e^(1/x) at infinity e^(1/infinity) = e^0 = 1

OpenStudy (anonymous):

Thanks!